OpenStudy (anonymous):

For sin(theta) = -0.4321, 0º < theta < 360º How many solutions are possible? In which quadrants would you find the solutions? Determine the reference angle for this equation to the nearest degree. Determine all the solutions to the equation to the nearest degree

OpenStudy (dumbcow):

There are 2 solutions - If you look at graph of sin, you notice it goes down from 0 to -1 then back up to 0, therefore it crosses value -.43 twice 3rd and 4th quadrant - see above arcsin(-0.4321) = -25.6 degrees = 360-25.6 = 334.4 since this is greater than 270 it is in 4th quadrant

OpenStudy (anonymous):

this looks like a great solution :)

OpenStudy (anonymous):

is the part below the answer to d? "arcsin(-0.4321) = -25.6 degrees = 360-25.6 = 334.4 since this is greater than 270 it is in 4th quadrant"

OpenStudy (dumbcow):

yes its half the answer now to find the solution in 3rd quadrant add 25.6 to 180

OpenStudy (anonymous):

study the sine graph

OpenStudy (anonymous):

the reference angle is -25.6

OpenStudy (anonymous):

thanks oodles!